# Energy functions of general dimensional diamond crystals based on the Kitaev model

### v3

@@ 3. The Kitaev model
$\sigma^{x}=\sqrt{-1}c_{1}c_{4},\ \ \sigma^{y}=\sqrt{-1}c_{2}c_{4},\ \ \sigma^{z}=\sqrt{-1}c_{3}c_{4}.$
@@
$\sigma^{x}_{v}=\sqrt{-1}c_{1}^{v}c_{4}^{v},\ \ \sigma^{y}_{v}=\sqrt{-1}c_{2}^{v}c_{4}^{v},\ \ \sigma^{z}_{v}=\sqrt{-1}c_{3}^{v}c_{4}^{v}.$
@@ 6. Fourier transform
In this section, for the energy function $\xi(\bm{q})$, we discribe zeros and energy gaps. The corresponding results in the case $d=2$ are due to A.Kitaev [4].
@@
We suppose that there exists a $(d+1)$-gon such that the lengths of the sides are $a_{0},a_{1},\cdots,a_{d}$. Since a side is the shortest length connecting two endpoints of a edge of a polygon, the inequalities (7.2) holds.
@@
First, we consider the case $d=2$. Then the statement holds becouse of the triangle inequality.
@@
In both cases, by hypothesis of induction there exists a $d$-gon such that the lengths of the sides are $e,a_{1},\cdots,a_{d-1}$. We attach the $d$-gon and the triangle by identifying them along the side of the length $e$. This construction yields a $(d+1)$-gon. By choosing $\epsilon$ sufficiently small, the two polygons sharing the side of length $e$ can be arranged so that they do not overlap. Therefore there exists a $(d+1)$-gon such that the lengths of the sides are $a_{0},a_{1},\cdots,a_{d}$.
@@ 8. Relation with the tight binding model
Let $G$ be the crystallographic group for $\Delta_{d}$. We suppose that $V(x)$ is invariant by the action of $G$. This Schrödinger operator $\hat{H}$ is a linear map $L^{2}(\check{D})\to L^{2}(\check{D})$ where $\check{D}$ is a fundamental domain of the reciprocal lattice of $A_{d}$ in $W^{*}$.
As an approximation model for the Schrödinger equation, we define the tight binding model for $\Delta_{d}$ by using the base graph $X_{0}$. We consider a family of functions $\psi_{v}\in L^{2}(\check{D})$, $v\in V(\Delta_{d})$ satisfying the following condtions
@@
    $\langle\psi_{v}|\psi_{v^{\prime}}\rangle=\int_{\check{D}}\psi_{v}^{*}\psi_{v^{\prime}}d\bm{x}=\delta_{v,v^{\prime}}$ for $v,v^{\prime}\in V(\Delta_{d})$.
@@
$h_{v,v^{\prime}}=\int_{\check{D}}{\psi_{v}}^{*}\hat{H}\psi_{v^{\prime}}dx$

### v4

@@ 3. The Kitaev model
$\sigma^{x}=-\sqrt{-1}c_{1}c_{4},\ \ \sigma^{y}=\sqrt{-1}c_{2}c_{4},\ \ \sigma^{z}=-\sqrt{-1}c_{3}c_{4}.$
@@
$\sigma^{x}_{v}=-\sqrt{-1}c_{1}^{v}c_{4}^{v},\ \ \sigma^{y}_{v}=\sqrt{-1}c_{2}^{v}c_{4}^{v},\ \ \sigma^{z}_{v}=-\sqrt{-1}c_{3}^{v}c_{4}^{v}.$
@@ 6. Fourier transform
In this section, for the energy function $\xi(\bm{q})$, we describe zeros and energy gaps. The corresponding results in the case $d=2$ are due to A.Kitaev [4].
@@
We suppose that there exists a $(d+1)$-gon such that the lengths of the sides are $a_{0},a_{1},\cdots,a_{d}$. Since a side is the shortest length connecting two endpoints of a edge of a polygon, the inequalities (7.2) hold.
@@
First, we consider the case $d=2$. Then the statement holds because of the triangle inequality.
@@
In both cases, by hypothesis of induction there exists a $d$-gon such that the lengths of the sides are $e,a_{1},\cdots,a_{d-1}$. We attach the $d$-gon and the triangle by identifying them along the side of the length $e$. This construction yields a $(d+1)$-gon. By choosing $\epsilon$ sufficiently small, the two polygons sharing the side of length $e$ can be arranged so that they do not overlap. Therefore there exists a $(d+1)$-gon such that the lengths of the sides are $a_{0},a_{1},\cdots,a_{d}$. ∎
@@ 8. Relation with the tight binding model
Let $G$ be the crystallographic group for $\Delta_{d}$. We suppose that $V(x)$ is invariant by the action of $G$. This Schrödinger operator $\hat{H}$ is a linear map $L^{2}(\mathbb{R}^{d})\to L^{2}(\mathbb{R}^{d})$. As in section 4 and 5, the domain $D^{\prime}_{\beta_{j}}$ is a fundamental domain of $A_{d}$ in $W$ and the domain $\check{D}$ is a fundamental domain of the reciprocal lattice of $A_{d}$ in $W^{*}$. Then, the operator $\hat{H}$ on $L^{2}(\mathbb{R}^{d})$ is unitarily equivalent to the direct-integral decompositions of the Hamiltonians $\hat{H}(\bm{q})$ acting on $L^{2}(D^{\prime}_{\beta_{j}})$ with respect to $\check{D}$ (see, for example in P. Kuchment [7], section 4).
As an approximation model for the Schrödinger equation, we define the tight binding model for $\Delta_{d}$ by using the base graph $X_{0}$. The space $L^{2}_{loc}(\mathbb{R}^{d})$ denotes
$\left\{f:\mathbb{R}^{d}\to\mathbb{C}\;\middle|\;f\in L^{2}(V)\text{ for every compact }V\subset\mathbb{R}^{d}\right\}.$
We consider a family of functions $\psi_{v}\in L^{2}_{loc}(\mathbb{R}^{d})$, $v\in V(\Delta_{d})$ satisfying the following conditions.
@@
    For each fixed $\bm{q}\in\check{D}$, $\langle\psi_{v}|\psi_{v^{\prime}}\rangle=\int_{D^{\prime}_{\beta_{j}}}\psi_{v}^{*}\psi_{v^{\prime}}d\bm{x}=\delta_{v,v^{\prime}}$ for $v,v^{\prime}\in V(\Delta_{d})$.
@@
$h_{v,v^{\prime}}=\int_{D^{\prime}_{\beta_{j}}}{\psi_{v}}^{*}\hat{H}\psi_{v^{\prime}}dx$